# Completely regular growth solutions to linear differential equations with exponential polynomials coefficients

### v1

@@ Abstract
Consider the linear differential equation
$f^{(n)}+A_{n-1}f^{(n-1)}+\cdots+A_{0}f=0$ (+)( + )
where the coefficients $A_{j},j=0,\ldots,n-1,$ are exponential polynomials. It is known that every solution is entire. This paper will show that all transcendental solutions of finite growth order are of completely regular growth. This problem was raised in Heittokangas et al.[8, p. 33], which involves an extensive question about Gol’dberg-Ostrovskiǐ’s Problem [5, p. 300]. Moreover, we define functions in a generalized class concluding exponential polynomial functions, which are also of completely regular growth.
Keywords: completely regular growth, exponential polynomials, Stokes rays, matrix asymptotic solutions
2020 MSC: 30D15, 30D35
@@ 1 Introduction
An exponential polynomial is an entire function of the form
$G=P_{1}e^{Q_{1}}+\cdots+P_{l}e^{Q_{l}}$ (1.1)
where $P_{j},Q_{j},j=1,\ldots,l,$ are polynomials, $l\in\mathbb{N}$. The Phragmén-Lindelöf indicator function of an entire function $f(z)$ of finite order $\rho(f)>0$ is given by
@@
for $\theta\in[0,2\pi)$. We say that $f$ is of completely regular growth (briefly denoted by c.r.g., see [9, pp. 139-140] or [11, pp. 6-8]) if there exists a sequence of Euclidean discs $D(z_{k},r_{k})$ satisfying
@@
as $r\to\infty$ uniformly in $\theta$.
Detailed studies of functions of c.r.g. can be found in [9] and [11]. Transcendental exponential polynomial functions are of completely regular growth, see [6, Lemma 1.3]. Petrenko [10, pp. 104-112] and Steinmetz [12] independently proved that the solutions of linear differential equations with polynomial coefficients are of c.r.g.. Gol’dberg and Ostrovskiǐ stated a classic problem in the book [5, p. 300]:
(Gol’dberg-Ostrovskiǐ) Suppose that $f$ is a finite order entire transcendental solution of
$f^{(n)}+a_{n-1}f^{(n-1)}+\cdots+a_{1}f^{\prime}+a_{0}f=0,$
whose coefficients are entire functions of completely regular growth. Is $f$ also of completely regular growth?
Recently Bergweiler [1] has constructed a counter-example about G.-O.’s problem. Therefore, Heittokangas et al.[8, p. 33] propose a relevant question by strengthening the assumption.
If the coefficients $a_{0}(z),\ldots,a_{n-1}(z)$ of
@@
Our main work is to positively answer the problem above. We give an estimation of $\log|f|$ on every small sector except for some areas around a finite number of rays from the origin, resulting $f$ of c.r.g., see Theorem 3.2 as well as Theorem 3.6.
Here is the structure of the paper. Section 2.1 is the brief introduction of c.r.g. on the rays with a useful lemma; In Section 2.2, we present Steinmetz’s result for $\log|f|$ regarding solutions $f$ of linear differential equations with rational function coefficients in the whole complex plane, and then we use techniques from Wasow’s book [14] to modify it to expressions for $\log|f|$ regarding solutions $f$ of linear differential equations with asymptotically expressed coefficients in some sector. Contents of Section 3 cover the main results and corresponding proofs of Theorems in Section 3 are stated in Section 4; In Section 5 we discuss some generalizations. There we define a new class of functions, generalizing the class of exponential polynomials in the class of functions of completely regular growth.
@@ 2 Basic concepts and lemmas
2.1 Compactification of the angular set for functions of completely regular growth on the rays
@@
where $E_{r}$ denotes the intersection of the set $E$ with the interval $(0,r)$, and mes$(E_{r})$ is the measure of $E_{r}$. It is obvious that the set of those values of $r$ for which the circle $|z|=r$ intersects a $C_{0}$-set is of zero relative measure. Refering to the definitions in [9, p. 139], a function $f$ that is holomorphic in $\{re^{i\theta}:\theta_{1}<\theta<\theta_{2},\theta_{1},\theta_{2}\in\mathbb{R},r>0\}$ and of order $\rho$ will be called a function of completely regular growth on the ray $\arg z=\theta$ for some $\theta\in(\theta_{1},\theta_{2})$ if the limit
@@
exists, where $E_{\theta}$ is a set of zero relative measure depending on $\theta$. Denote $R_{\theta}=\{re^{i\theta},r\notin E_{\theta},r>0\}$ for some angle $\theta$, a subset $\mathbb{M}\subset(\theta_{1},\theta_{2})$ and
@@
$H_{f}(r,\theta)=\frac{\log|f(re^{i\theta})|}{r^{\rho}},r\notin\displaystyle\bigcup_{\theta\in\mathbb{M}}E_{\theta},$
converges uniformly to $h_{f}(\theta)$ for all $\theta,\theta\in\mathbb{M}.$ Additionally, Levin [9, p. 140] has proved that $f(z)$ is a function of completely regular growth on every ray which is a limiting ray of the set of rays along which $f(z)$ is of completely regular growth.
@@ 2.2 The asymptotic solutions of differential equations in a sector
with entire coefficients $a_{0}(z),\ldots,a_{n-1}(z)$, Petrenko [10, pp. 104–112] has shown that transcendental solutions to linear differential equations (2.1) with polynomial coefficients are of completely regular growth. Steinmetz[12, Theorem 1] has proved a more generalized theorem in the following. The differential equation
@@
allows rational coefficients $R_{0}(z),\ldots,R_{n-1}(z)$. And it is always assumed the solutions of (2.2) are meromorphic functions.
@@
This kind of exceptional rays are called Stokes rays of $f(z).$ Concerning the definition of Stokes ray, one can refer to Steinmetz[12], and the general definition in Wasow[14, Sec. 15].
In proving our results, our goal is not to assume anymore that the coefficients of the equation (2.2) are rational functions, but to assume that the coefficients have an asymptotic expansion at $\infty$ in a sector $S$. From this point of view, it is necessary to define the asymptotic symbol $\sim$. We write
$f(x)\sim C_{0}+\frac{C_{1}}{x}+\frac{C_{2}}{x^{2}}+\cdots,$
@@
$\lim_{x\to\infty}\left(f(x)-C_{0}-\frac{C_{1}}{x}-\frac{C_{2}}{x^{2}}-\cdots-\frac{C_{m}}{x^{m}}\right)x^{m}=0,$
@@
$f(x)=C_{0}+\frac{C_{1}}{x}+\frac{C_{2}}{x^{2}}+\cdots+\frac{C_{m}}{x^{m}}+O\left(\frac{1}{x^{m+1}}\right).$
@@
$z^{h}\frac{d\bm{Y}}{dz}=\bm{B}(z)\bm{Y},\enspace h\in\mathbb{N},$ (2.4)
where $\bm{Y},\bm{B}$ are $n$-by-$n$ matrixes, $n\in\mathbb{N^{*}}$, and the matrix $\bm{B}(z)$ is holomorphic at $z=0$. When $h>1,$ the singular point is irregular. We want to make a transformation $x=1/z$ taking into the equation (2.4), affording
$x^{-d}\frac{d\bm{Y}}{dx}=\bm{A}(x)\bm{Y},$ (2.5)
with $d=h-2$ and $\bm{A}(x)=-\bm{B}(1/z)$. The integer $d+1$ is called the rank of the singular point. Then for $n$-by-$n$ constant matrixes $\bm{A}_{r}$, the matrix $\bm{A}(x)$ could have its expansion with form
$\bm{A}(x)=\sum_{j=0}^{\infty}\bm{A}_{j}x^{-j},\quad|x|>x_{0},$ (2.6)
which is holomorphic at $x=\infty$, or an expansion of this form asymptotic to $\bm{A}(x)$ in some sector $S$, i.e.
$\bm{A}(x)\sim\sum_{j=0}^{\infty}\bm{A}_{j}x^{-j},\enspace x\in S.$ (2.7)
We pay more interest to the latter case (2.7) and wonder if there exists an asymptotic expression of the solution $\bm{Y}$ in the same area. The matrix solution in the sector $S$ is summarized in the following theorem, see [14, p.111]:
@@
Let $\bm{A}(x)$ be an $n$-by-$n$ matrix function, which has the asymptotic form (2.7) in $S$. Then, the differential equation (2.5) possesses a fundamental matrix solution of the form
$\bm{Y}(x)=\hat{\bm{Y}}(x)x^{\bm{G}}e^{\bm{Q}(x)}$ (2.8)
corresponding to every sufficiently narrow subsector of $S$. Here $\hat{\bm{Y}}(x)$ permits an asymptotic series in power of $x^{-1/p},p\in\mathbb{N^{*}}$, in this subsector as $x\rightarrow\infty;$ $\bm{G}$ is a constant matrix, $\bm{Q}(x)$ is a diagonal matrix whose diagonal elements are polynomials in $x^{1/p}$.
@@
$f^{(n)}(x)+A_{n-1}(x)f^{(n-1)}(x)+\cdots+A_{1}(x)f^{\prime}(x)+A_{0}(x)f(x)=0$ (2.9)
possesses coefficients $A_{k}(x)$ with the asymptotic expansions
$A_{k}(x)\sim x^{d_{k}}\sum_{j=0}^{\infty}A_{kj}x^{-j},\enspace x\in S,A_{kj}\in\mathbb{C},d_{k}\in\mathbb{N},k=0,\cdots,n-1.$
@@
$x^{-d}\frac{d\bm{y}}{dx}=\bm{A}(x)\bm{y},\enspace x\in S,$
with $x^{d}\bm{A}=\tilde{\bm{A}}$ and integer $d=\max\limits_{k=0,\ldots,n-1}\{d_{k}\}$. Therefore, we transform equation (2.9) to a linear system (2.5) with $\bm{A}$ in the form of (2.7) after set $\bm{Y}=\{\bm{y}_{1},\ldots,\bm{y}_{n}\}$, $\bm{y}_{i}=\{f_{i},\ldots,f^{(n-1)}_{i}\}^{T},$ where $f_{i}$ are $n$ linearly independent solutions of (2.9), $i=1,\ldots,n$. The elements of $\bm{A}_{j}$ in (2.7) consisted of some of the $A_{kj},0,1$, and $k=0,\ldots,n-1,j\in\mathbb{N}.$
By Lemma 2.3 and fundamental calculations of elements of Matrixes, the $n$ linear independent solutions $f_{i}$ in $x$ in $S$ are
$f_{i}(x)=e^{\varLambda_{i}(x^{\frac{1}{p}})}x^{c_{i}}\varOmega_{i}(x,\log x),\quad 1\leq i\leq n.$ (2.10)
Here $\varLambda_{i}$ is a polynomial in $x^{1/p}$, $p\in\mathbb{N^{*}}$, $c_{i}\in\mathbb{C}$, and $\varOmega_{i}$ is a polynomial in $\log x$ with coefficients $\beta_{i,j}(x)$ possessing asymptotic forms
$\beta_{i,j}(x)\sim\sum_{s=0}^{\infty}\alpha_{i,j,s}x^{-s/p},\quad\alpha_{s}\in\mathbb{C},1\leq i\leq n,0\leq j\leq n-1.$
@@
Let $f(x)$ be any solution of (2.9) in the sector $S$ and let $\arg z=\theta$ be an arbitrary direction in the interior of $S$. Then there is $p\in\mathbb{N},h>0,$ a polynomial $Q_{\theta}$ in $z^{1/p}$, such that
$\log|f(x)|=\Re Q_{\theta}(x^{1/p})+O(\log|x|)$ (2.11)
as $x\rightarrow\infty$ in $\theta\leq\arg x\leq\theta+h$ possibly outside an exceptional set consisting of
@@
    (countably many) disks $|x-x_{\mu}|<|x_{\mu}|^{1-\varepsilon},$ where $\varepsilon$ is positive and the counting function of the sequence {$x_{\mu}$} is $O(\log r)$;
@@
    a logarithmic semi-strip $0\leq\arg x-\theta<\dfrac{\log^{+}|x|}{|x|^{\frac{1}{p}}}$.
@@
Note that it is always possible to take the small sector in $S$ for $\theta\leq\arg x\leq\theta+h,h>0,$ when the ray $\arg z=\theta$ is in the interior of $S$. The polynomial $Q_{\theta}$ depending on $\theta$, but is piecewise constant and changes only when $\arg z=\theta$ is a Stokes ray.
3 Main results

### v2

# On the Goldberg-Ostrovskii Problem for Linear Differential Equations with Exponential Polynomial Coefficients

@@ Abstract
The Goldberg–Ostrovskii problem asks whether finite-order solutions of a linear differential equation inherit the property of completely regular growth (c.r.g.) from its coefficients. While Bergweiler’s counterexample demonstrated that the answer is negative in general, this paper proves that when the coefficients are restricted to the classical and rich subclass of exponential polynomials, the regularity transmission does hold. Thereby we affirm the conjecture posed by Heittokangas, Ishizaki, Tohge and Wen.
Our results reveal the closed nature of exponential polynomials in the context of regularity transfer from the perspective of equation dynamics, and provide a new perspective for the study of the structure of related function classes.
Keywords: Exponential polynomials; Completely regular growth; Goldberg-Ostrovskii’s problem; Asymptotic analysis
2020 MSC: 30D15, 34M10, 30D35
@@ 1 Introduction
Completely regular growth and classic problems on equations
The theory of linear differential equations in the complex domain represents one of the most profound intersections of complex analysis, operator theory, and dynamical systems. At the heart of this theory lies the fundamental problem of understanding the growth and distribution of zeros of entire function solutions. The Ukrainian mathematician B. Ja. Levin revealed a crucial connection through the concept of completely regular growth (c.r.g.), which establishes a deep relationship between growth control and zero distribution on sectors.
The Phragmén-Lindelöf indicator function of an entire function $f(z)$ of finite order $\rho(f)>0$ is given by
@@
for $\theta\in[0,2\pi)$. We say that $f$ is of completely regular growth (briefly denoted by c.r.g., see [23, pp. 139-140] or [32, pp. 6-8]) if there exists a sequence of Euclidean discs $D(z_{k},r_{k})$ satisfying
@@
as $r\to\infty$ uniformly in $\theta$. A function of c.r.g. is equivalent to the regular distribution of its zeros, see [23, p. 158].
A. A. Goldberg and I. V. Ostrovskii stated a classic problem in the book [15, p. 300]:
(Goldberg-Ostrovskii) Suppose that $f$ is a finite order transcendental solution of f^(n)+a_n-1f^(n-1)++a_1f’+a_0f=0, whose coefficients are entire functions of c.r.g.. Is $f$ also of c.r.g.?
This problem was formulated by V.P. Petrenko in [29, p. 132] but he omitted the condition of c.r.g. of the coefficients $a_{j}(z)$, for $1\leq j\leq n-1$. As the book was published posthumously without the author’s final proofreading, this omission may have been accidental. Crucially, without this c.r.g. condition, the answer to the problem is negative, see [15, p. 300]. Bergweiler [4] constructed a counterexample showing that additional conditions are necessary, thus motivating the search for the precise conditions under which regularity is inherited. This makes us wonder, what should the “missed condition” on coefficients be that makes all the finite order solutions of c.r.g.?
We wish to recognize just what functions are of c.r.g., which is an important consideration on recapturing this problem. Detailed studies of functions of c.r.g. can be found in [23] and [32]. Obviously, non-transcendental functions can not be functions of c.r.g. and $e^{z}$ should be the simplest one without zeros.
The combination of exponential terms and polynomials, such as $1-z+e^{z}$, can be also easily verified from the condition (1.2). Mathematicians such as Polyà [30], Schwengeler [33] and Dickson [6] have found that the functions – exponential sums, see (3.1), has a distribution of zeros that asymptotic to some rays, i.e., just as $1-z+e^{z}$ does. Although they give different estimates of the asymptotic trend, the regularly distributed zero property tells us such functions are equivalently of c.r.g..
A more general function class comparing exponential sums is known to us below. An exponential polynomial is an entire function of the form
$G=P_{1}e^{Q_{1}}+\cdots+P_{l}e^{Q_{l}}$ (1.3)
where $P_{j},Q_{j},j=1,\ldots,l,$ are polynomials, $l\in\mathbb{N}$. It is known that exponential sums are exponential polynomials with all $Q_{j}$ of degree 1. Transcendental exponential polynomial functions are also of c.r.g., see [17, Lemma 1.3] for example.
For the more complicated functions of c.r.g., it was not possible to write the explicit forms, but it was found that the exponential terms played a crucial role. To explain it clearly, we know different exponential terms control the growth of an exponential sum in different sectors. Such as $e^{-z}$ controls the growth of $e^{-z}+e^{z}$ on the left-half plane and $e^{z}$ controls it on the right-half plane. In the area around the positive and negative imaginary axes, there will be some sudden changes in the function values to produce zeros, such a phenomenon is known as the Stokes phenomenon, see Wasow [38, Sec. 15].
Although the asymptotic analysis theory on Stokes pheonomenon does not only deals with entire functions, such as also the example – Hankel functions – which are the linearly independent solutions of Bessel’s differential equation introduced in [38, p. 78], the well-known Airy functions $\mathrm{Ai}(z)$ and $\mathrm{Bi}(z)$, as the solutions of the Airy equation
$f^{\prime\prime}-zf=0,$
has the typical Stokes phenomenon on the negative axis, see [27, Chap. 9]. Petrenko [29, pp. 104-112] and Steinmetz [34] independently proved that the entire solutions of linear differential equations with polynomial coefficients (Steinmetz indicated this result also holds for rational coefficients) are of c.r.g.. More such entire special functions of c.r.g., e.g. generalized Airy functions $A_{n}(z)$ and $B_{n}(z)(n\in\mathbb{Z})$, Bessel functions of the first kind $J_{n}(z)(n\in\mathbb{Z})$, and Kummer functions $M(a,b,z)(a,b\in\mathbb{C})$, can be seen separately in [27, §9.13, §10.2, §13.2].
The fact that so many c.r.g. special functions (or associated forms) fit into the framework of exponential polynomials indicates that this class is a natural candidate for the study of regularity transmission—being both broad yet well-behaved. Any exponential polynomial can still be constructed to be a solution to a linear differential equation whose coefficients are polynomials, see [37]. Heittokangas et al. are expert in exponential polynomials and they indicated the regularity that for certain LDE the class of finite order solutions is always larger (at least not smaller) than the class of coefficients, see the survey [19] on exponential polynomials. Therefore, in [19, p. 33] they proposed a relevant question by strengthening the assumption on G.-O.’s problem.
(H.-I.-T.-W.) If the coefficients $a_{0}(z),\ldots,a_{n-1}(z)$ of
@@
Our recent work in [24] provides a partial answer to this line of inquiry, showing that for linear differential equations whose coefficients are exponential sums (a subclass of exponential polynomials) with rational frequencies, all finite-order solutions are themselves exponential sums and consequently of completely regular growth.
This paper is to give a complete answer to this problem. We give an estimation of $\log|f|$ on every small sector except for some areas around a finite number of rays from the origin, resulting $f$ of c.r.g.. See Theorem 3.2, Theorem 3.5, and Theorem 3.9. In summary, this paper completely proves that every finite-order transcendental entire solution of a linear differential equation with exponential polynomial coefficients is of completely regular growth. This conclusively settles the Heittokangas–Ishizaki–Tohge–Wen conjecture.
Organization
Section 2 recalls necessary concepts: functions of c.r.g. on rays and its compactification, the asymptotic theory of ODEs in sectors by Wasow, and Steinmetz’s application to LDEs. Section 3 presents our main results for equations with exponential sum coefficients (Theorem 3.2, Theorem 3.5) and with exponential polynomial coefficients (Theorem 3.9), including a description of the asymptotic sectors and exceptional sets. Section 4 contains the complete proofs, detailing the reduction to polynomial-coefficient equations in sectors, the analysis near critical and Stokes rays, and the final verification of the c.r.g. property via the compactness principle. Section 5 offers concluding remarks and outlines future directions, including the exploration of the solution space structure via operator-theoretic methods. Section 6 (Appendix) provides a necessary modification of Wasow’s sectorial asymptotic theorem to regions with holes, which underpins the proof when dealing with exceptional sets.
@@ 2 Basic concepts and lemmas
2.1 Compactification for functions of c.r.g. on the rays
@@
[…]
In the definition of c.r.g., the exception condition $z\notin C_{0}$ and $|z|\notin E_{0}$ is equivalent by the reason of the correspondence of regular distribution of zeros, see [23, Thm. 1&2, Chap. 1] and [23, Thm.4, p. 158], or [32, Thm. 1.2.1].
Refering to the definitions in [23, p. 139], a function $f$ that is holomorphic in $\{re^{i\theta}:\theta_{1}<\theta<\theta_{2},\theta_{1},\theta_{2}\in\mathbb{R},r>0\}$ and of order $\rho$ will be called a function of completely regular growth on the ray $\arg z=\theta$ for some $\theta\in(\theta_{1},\theta_{2})$ if the limit
@@
exists, where $E_{\theta}$ is a $E_{0}$-set depending on $\theta$. Denote $R_{\theta}=\{re^{i\theta},r\notin E_{\theta},r>0\}$ for some angle $\theta$, a subset $\mathbb{M}\subset(\theta_{1},\theta_{2})$ and
@@
$H_{f}(r,\theta)=\frac{\log|f(re^{i\theta})|}{r^{\rho}},~r\notin\displaystyle\bigcup_{\theta\in\mathbb{M}}E_{\theta},$
converges uniformly to $h_{f}(\theta)$ for all $\theta,\theta\in\mathbb{M}.$ Additionally, Levin [23, p. 140] has proved that $f(z)$ is a function of completely regular growth on every ray which is a limiting ray of the set of rays along which $f(z)$ is of completely regular growth.
@@ 2.2 The asymptotic solutions of differential equations in a sector
with entire coefficients $a_{0}(z),\ldots,a_{n-1}(z)$, Petrenko [29, pp. 104–112] has shown that transcendental solutions to linear differential equations (2.1) with polynomial coefficients are of completely regular growth. Steinmetz [34, Theorem 1] has proved a more generalized theorem in the following. The differential equation
@@
allows rational coefficients $R_{0}(z),\ldots,R_{n-1}(z)$. And it is only considered the solutions of (2.2) are meromorphic functions.
@@
This kind of exceptional rays are called Stokes rays of $f(z).$ Concerning the definition of Stokes ray, one can refer to Steinmetz [34], and the general definition in Wasow [38, Sec. 15]. In proving our results, our goal is not to assume anymore that the coefficients of the equation (2.2) are rational functions, but to assume that the coefficients have an asymptotic expansion at $\infty$ in a sector $S$. From this point of view, it is necessary to define the asymptotic symbol $\sim$. We write
$f(z)\sim C_{0}+\frac{C_{1}}{z}+\frac{C_{2}}{z^{2}}+\cdots,$
@@
$\lim_{z\to\infty}\left(f(z)-C_{0}-\frac{C_{1}}{z}-\frac{C_{2}}{z^{2}}-\cdots-\frac{C_{m}}{z^{m}}\right)z^{m}=0,$
@@
$f(z)=C_{0}+\frac{C_{1}}{z}+\frac{C_{2}}{z^{2}}+\cdots+\frac{C_{m}}{z^{m}}+O\left(\frac{1}{z^{m+1}}\right).$
@@
$x^{h}\frac{d\bm{Y}}{dx}=\bm{B}(x)\bm{Y},\kern 5.0pth\in\mathbb{N},$ (2.4)
where $\bm{Y},\bm{B}$ are $n$-by-$n$ matrixes, $n\in\mathbb{N^{*}}$, and the matrix $\bm{B}(x)$ is holomorphic at $x=0$. When $h>1,$ the singular point is irregular. We want to make a transformation $z=1/x$ taking into the equation (2.4), affording
$z^{-d}\frac{d\bm{Y}}{dz}=\bm{A}(z)\bm{Y},$ (2.5)
with d=h2d=h-2 and $\bm{A}(z)=-\bm{B}(1/x)$. The integer $d+1$ is called the rank of the singular point. Then for $n$-by-$n$ constant matrixes $\bm{A}_{r}$, the matrix $\bm{A}(z)$ could have its expansion with form
$\bm{A}(z)=\sum_{j=0}^{\infty}\bm{A}_{j}z^{-j},\quad|z|>z_{0},$ (2.6)
which is holomorphic at $z=\infty$, or an expansion of this form asymptotic to $\bm{A}(z)$ in some sector $S$, i.e.
$\bm{A}(z)\sim\sum_{j=0}^{\infty}\bm{A}_{j}z^{-j},\kern 5.0ptz\in S.$ (2.7)
We pay more interest to the latter case (2.7) and wonder if there exists an asymptotic expression of the solution $\bm{Y}$ in the same area. The matrix solution in the sector $S$ is summarized in the following theorem, see [38, p.111]:
@@
Let $\bm{A}(z)$ be an $n$-by-$n$ matrix function, which is holomorphic and has the asymptotic form (2.7) in $S$. Then, the differential equation (2.5) possesses a fundamental matrix solution of the form
$\bm{Y}(z)=\hat{\bm{Y}}(z)z^{\bm{G}}e^{\bm{Q}(z)}$ (2.8)
corresponding to every sufficiently narrow subsector of $S$. Here $\hat{\bm{Y}}(z)$ permits an asymptotic series in power of $z^{-1/p},p\in\mathbb{N^{*}}$, in this subsector as $z\rightarrow\infty;$ $\bm{G}$ is a constant matrix, $\bm{Q}(z)$ is a diagonal matrix whose diagonal elements are polynomials in $z^{1/p}$.
@@
possesses coefficients $A_{k}(z)$ with the asymptotic expansions
$A_{k}(z)\sim z^{d_{k}}\sum_{j=0}^{\infty}A_{kj}z^{-j},\kern 5.0ptz\in S,A_{kj}\in\mathbb{C},d_{k}\in\mathbb{N},k=0,\cdots,n-1.$
@@
$z^{-d}\frac{d\bm{y}}{dz}=\bm{A}(z)\bm{y},\kern 5.0ptz\in S,$
with $z^{d}\bm{A}=\tilde{\bm{A}}$ and integer $d=\max\limits_{k=0,\ldots,n-1}\{d_{k}\}$. Therefore, we transform equation (2.9) to a linear system (2.5) with $\bm{A}$ in the form of (2.7) after set $\bm{Y}=\{\bm{y}_{1},\ldots,\bm{y}_{n}\}$, $\bm{y}_{i}=\{f_{i},\ldots,f^{(n-1)}_{i}\}^{T},$ where $f_{i}$ are $n$ linearly independent solutions of (2.9), $i=1,\ldots,n$. The elements of $\bm{A}_{j}$ in (2.7) consisted of some of the $A_{kj},0,1$, and $k=0,\ldots,n-1,j\in\mathbb{N}.$
By Lemma 2.3 and fundamental calculations of elements of Matrixes, the $n$ linear independent solutions $f_{i}$ in $z$ in $S$ are
$f_{i}(z)=e^{\varLambda_{i}(z^{\frac{1}{p}})}z^{c_{i}}\varOmega_{i}(z,\log z),\quad 1\leq i\leq n.$ (2.10)
Here $\varLambda_{i}$ is a polynomial in $z^{1/p}$, $p\in\mathbb{N^{*}}$, $c_{i}\in\mathbb{C}$, and $\varOmega_{i}$ is a polynomial in $\log z$ with coefficients $\beta_{i,j}(z)$ possessing asymptotic forms
$\beta_{i,j}(z)\sim\sum_{s=0}^{\infty}\alpha_{i,j,s}z^{-s/p},\quad\alpha_{s}\in\mathbb{C},1\leq i\leq n,0\leq j\leq n-1.$
@@
Let $f(z)$ be any solution of (2.9) in the sector $S$ and let $\arg z=\theta$ be an arbitrary direction in the interior of $S$. Then there is $p\in\mathbb{N},h>0,$ a polynomial $Q_{\theta}$ in $z^{1/p}$, such that
$\log|f(z)|=\Re Q_{\theta}(z^{1/p})+O(\log|z|)$ (2.11)
as $z\rightarrow\infty$ in $\theta\leq\arg z\leq\theta+h$ possibly outside an exceptional set consisting of
@@
3 Main results on solutions of c.r.g.